OpenStudy (anonymous):

collect like terms 14x+7y-11x-9y =--

OpenStudy (anonymous):

its actually quite simple, you just need to look at it like this (14x-11x)+(7y-9y) = ---. you can rearrange it like that and you can group (use parenthesis) like i did when dealing with simple algebra. So I assume you know how to subtract 11 from 14 and 9 from 7? and then those differences will be the coefficients. respectively.

OpenStudy (anonymous):

urbanderivative would the answer be 6

OpenStudy (anonymous):

how did you come up with that answer. It's incorrect, but I need to know how to fix you.

OpenStudy (anonymous):

i meant 5

OpenStudy (anonymous):

close, but you have to know that x and y are different variables and you cannot subtract them.

OpenStudy (anonymous):

is it 19

OpenStudy (anonymous):

you are leaving out your variables. You answer must include both x and y.

OpenStudy (anonymous):

i don't understand pls show me

OpenStudy (anonymous):

(14x-11x)+(7y-9y) = --- where 14x-11x is 3x. That's simple subtraction right? so now we have (14x-11x)+(7y-9y) = 3x + (?)

OpenStudy (anonymous):

\[(14x-11x) = x(14-11)= x(3) = 3x\]

OpenStudy (anonymous):

so my answer is 3xcorrect